OpenStudy (anonymous):

Need help solving a recurrence definition in combinatorics:

OpenStudy (anonymous):

Passwords for a certain computer system are strings of uppercase letters. A valid password must contain an even number of X's. Determine a recurrence relation for the number of valid passwords of length n.

OpenStudy (anonymous):

adding an x to a bad password would work, adding any other letter to a bad password would work

OpenStudy (fwizbang):

let a_n be the no of good pass words with n letters and b_n be the number of bad passwords with n letters. a_n+b_n = (26)^n a_(n+1) = 25*a_n + b_n b_(n+1)= a_n +25 b_n ?

OpenStudy (anonymous):

is it legal to define this way, or are we only allowed one variable

OpenStudy (fwizbang):

you can always eliminate b_n using the first equation

OpenStudy (anonymous):

how would that look if you did it eliminating b_n